# Corrugation-dominated mechanical softening of defect-engineered graphene

### v1

@@ Acknowledgments
The graphene with its Poly(methyl methacrylate) (PMMA) sacrificial layer was put into de-ionized water floating on the water surface. Next, the material was scooped out by the TEM chip followed by subsequent annealing in air at 150C. Then, the PMMA sacrificial layer was dissolved in an acetone bath at $50^{\circ}$. The necessity for a custom design stems from the need of a perforated support required for STEM that is also mechanically sufficiently rigid for the normal forces applied during AFM nanoindentation. To the best of our knowledge, the latter is not fulfilled by most commercially available support chips due to lack of thickness of the SiN support membrane, as demonstrated in the Supplemental Material [28], where the divergence between the curves acquired on the rigid Si wafer frame and the SiN membrane is a result of membrane bending. Comparing the two shows drastically reduced membrane bending on the custom support with 1000 nm thickness and $90\times 90$ $\mu$m2 wide windows, which is sufficiently low for the forces applied during AFM nanoindentation. Another advantage of the custom design is that the same sample location can be found in each instrument, which is enabled by a binary marker system of the SiN support windows arranged in a $3\times 3$ array, where a missing hole represents a one and an existing hole represents a zero, as shown in the Supplemental Material [28].
@@
The MAADF detector covered a semi-angular range of 60-200 mrad and the convergence angle of the electron beam was ca. 35 mrad. For each sample, the selected areas in the order of $10^{4}$ nm2 were defined by four corner points to establish the sample height at all positions through a bilinear interpolation. The selected area was dissected into increments with a nominal size of $5\times 5$ nm2, which were semi-automatically imaged at atomic resolution. In order to prevent overlap of the imaged areas, an offset of one or two images was used. To minimize residual stage movement during image acquisition, a sleep time of two or three seconds between stage translation and start of the acquisition was implemented.
@@
A machine learning force field, trained via active learning for distorted and ruptured membranes, was developed for a total of $12,000$ MD steps over $12$ different membrane systems and a total of $585$ configurations evaluated by density functional theory (DFT) using the Perdew–Burke–Ernzerhof functional [59]. Periodic boundary conditions in all directions were applied. To prevent interactions along the surface normal of the graphene sheet a vacuum of approximately $6$ nm was introduced. A total of $60$ simulations per vacancy density were performed. Each simulation started with a perfectly flat graphene sheet and a temperature of $T=400$ K (to accelerate the built-up of corrugation) in an isobaric-isothermal ensemble (NPT) using a Langevin thermostat. The system was then equilibrated until the average corrugation no longer increased, which required $20$ ps in the case of pristine graphene and $40-60$ ps for defective graphene. This was followed by $0.3$ ps of damped MD to relax the system, in order to obtain a good starting point for the subsequent procedure. Subsequently, the membrane was pulled at $T=300$ K by incrementally enlarging the in-plane cell size. This included a total of $20$ strain steps, where in each step the in-plane cell size was increased by $0.05$%. Each step was simulated for $10$ ps, in order to ensure that the system has sufficient time to adapt to the increased volume. The standard deviation of the atomic heights distribution was determined at each time step. This distribution was sampled over all atoms and simulated systems. The two-dimensional elastic modulus was then obtained by fitting the energy-strain curve with

### v3

@@ Acknowledgments
Data availability
The data that support the findings of this article are openly available [53].
@@
The Easy Transfer graphene from Graphenea Inc. graphene was put into de-ionized water floating on the water surface. Next, the material was scooped out by custom-designed SiN TEM chips purchased from Silson Ltd followed by subsequent annealing in air at 150C. Then, the Poly(methyl methacrylate) sacrificial layer was dissolved in an acetone bath at $50^{\circ}$. The necessity for a custom design stems from the need of a perforated support required for STEM that is also mechanically sufficiently rigid for the normal forces applied during AFM nanoindentation. To the best of our knowledge, the latter is not fulfilled by most commercially available support chips due to lack of thickness of the SiN support membrane, as demonstrated in the Supplemental Material [27], where the divergence between the curves acquired on the rigid Si wafer frame and the SiN membrane is a result of membrane bending. Comparing the two shows drastically reduced membrane bending on the custom support with 1000 nm thickness and $90\times 90$ $\mu$m2 wide windows, which is sufficiently low for the forces applied during AFM nanoindentation. Another advantage of the custom design is that the same sample location can be found in each instrument, which is enabled by a binary marker system of the SiN support windows arranged in a $3\times 3$ array, where a missing hole represents a one and an existing hole represents a zero, as shown in the Supplemental Material [27].
@@
The MAADF detector covered a semi-angular range of 60-200 mrad and the convergence angle of the 60-keV electron beam was ca. 35 mrad. For each sample, the selected areas in the order of $10^{4}$ nm2 were defined by four corner points to establish the sample height at all positions through a bilinear interpolation. The selected area was dissected into increments with a nominal size of $5\times 5$ nm2, which were semi-automatically imaged at atomic resolution. In order to prevent overlap of the imaged areas, an offset of one or two images was used. To minimize residual stage movement during image acquisition, a sleep time of two or three seconds between stage translation and start of the acquisition was implemented.
@@
A machine learning force field, trained via active learning for distorted and ruptured membranes, was developed for a total of $12,000$ MD steps over $12$ different membrane systems and a total of $585$ configurations evaluated by density functional theory (DFT) using the Perdew–Burke–Ernzerhof functional [61]. MD simulations were performed using VASP [62] with monoclinic simulation cells, and periodic boundary conditions in all directions were applied. To prevent interactions along the surface normal of the graphene sheet a vacuum of approximately $6$ nm was introduced. A total of $60$ simulations per vacancy density were performed. Each simulation started with a perfectly flat graphene sheet and a temperature of $T=400$ K (to accelerate the built-up of corrugation) in an isobaric-isothermal ensemble (NPT) using a Langevin thermostat. The system was then equilibrated until the average corrugation no longer increased, which required $20$ ps in the case of pristine graphene and $40-60$ ps for defective graphene. This was followed by $0.3$ ps of damped MD to relax the system, in order to obtain a good starting point for the subsequent procedure. Subsequently, the membrane was pulled at $T=300$ K by incrementally enlarging the in-plane cell size. This included a total of $20$ strain steps, where in each step the in-plane cell size was increased by $0.05$%. Each step was simulated for $10$ ps, in order to ensure that the system has sufficient time to adapt to the increased volume. The standard deviation of the atomic heights distribution was determined at each time step. This distribution was sampled over all atoms and simulated systems. The two-dimensional elastic modulus was then obtained by fitting the energy-strain curve with
@@
FIG. S 6: Simulated graphene membranes. Visualization of the simulated structures of pristine graphene, graphene with single vacancies and graphene with double vacancies, each containing up to $6,000$ atoms. The pristine case includes a schematic illustration of the viewing region shown in Fig. 4(b). The visualizations have been created by the VESTA software [47].